OpenStudy (anonymous):

In a simple random sample of 1000 people taken from Town A, 13% are senior citizens. In an independent simple random sample of 600 people taken from Town B, 17% are senior citizens. Is the percent of senior citizens different in the two towns? Or is this just chance variation? Choose the closest estimated standard error of the difference between the percents of senior citizens in the two samples, Under the null hypothesis. 1 or 1.2 or 1.4 or 1.6 or 1.8 or 2?

OpenStudy (amistre64):

\[z=\frac{mean1 - mean2}{\sqrt{\frac{s^2}{n1}+\frac{s^2}{n2}}}\] this sounds familiar ...

OpenStudy (amistre64):

the under part might be the standard error if im remembering it right

OpenStudy (amistre64):

can you recall what that might be from? http://www.wolframalpha.com/input/?i=sqrt%28%28.13*.87%2F1000%29%2B%28.17*.83%2F600%29%29

OpenStudy (anonymous):

is not like this?

OpenStudy (anonymous):

so is 1.8

OpenStudy (anonymous):

agree?

OpenStudy (amistre64):

standard error of the means ... which relates to proportions of course

OpenStudy (anonymous):

so do you agree?

OpenStudy (amistre64):

id have to review some stuff first :)

OpenStudy (anonymous):

Its correct do not waste your time i checked it ;) Thank you

OpenStudy (amistre64):

standard error of the mean:\[\frac{\sigma}{\sqrt{n}}\] using proportions that would amount to::\[\sqrt{\frac{pq}{n}}\] the difference between means tends to have an addition under there; and prolly by luck i got 1.8... as well

OpenStudy (anonymous):

the easiest way is always the best ;)